40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14523     Entry time: Mon Apr 8 18:28:25 2019     In reply to: 14521     Reply to this: 14524
Author: gautam 
Type: Update 
Category: ALS 
Subject: EX Green PDH checkout 

I worked on characterizing the green PDH setup at EX, as part of the ALS noise budgeting process. Summary of my findings:

  1. Green doubling efficiency is ~ 1.5 %/W (3mW of green for 450mW of IR). This is ~half of what was measured on the PSL table. There are probably large errors associated with power measurement with the Ophir power meter, but still, seems like a big mismatch.
  2. The green REFL photodiode is a Thorlabs PDA36A
    • It is being run on 30 dB gain setting, corresponding to a transimpedance of 47.5 kohm into high impedance loads. However, the PD bandwidth for this gain setting is 260 kHz according to the manual, whereas the PDH modulation sidebands on the green light are at twice the modulation frequency, i.e. ~560 kHz, so this is not ideal.
    • There was ~250 uW of green light incident on this photodiode, as measured with the Ophir power meter.
    • The DC voltage level was measured to be ~2.7 V on a scope (High-Z), which works out to ~280 uW of power, so the measurements are consistent.
    • When the cavity is locked, there is about 25% of this light incident on the PD, giving a shot noise level of ~25 nV/rtHz. The dark noise level is a little higher, at 40nV/rtHz.
    • Beam centering on the PD looked pretty good to the eye (it is a large-ish active area, ~3mmx3mm).
    • The beam does not look Gaussian at all - there are some kind of fringes visible in the vertical direction in a kind of halo around the main cavity reflection. Not sure what the noise implications of this are. I tried to capture this in a photo, see Attachment #1. Should an Iris/aperture be used to cut out some of this junk light before the reflection photodiode?
  3. The in-going beam was getting clipped on the Faraday Isolator aperture (it was low in pitch).
    • I fixed this by adjusting the upstream steering, and then moving the two PZT mounted green steering mirrors to recover good alignment to the X arm cavity.
    • GTRX level of ~0.5 was recovered.
  4. To estimate the mode-matching of the input beam to the cavity axis, I looked at the reflected light with the cavity locked, and with just the prompt reflection from the ETM:
    • DC light level on the reflection photodiode was monitored using the High-Z input o'scope.
    • Measured numbers are Plocked ~ 660 mV, Pmisaligned ~ 2.6V, giving a ratio of 0.253.
    • While locked, there was a ~ 10 Hz periodic variation in the DC light level on the green REFL photodiode - not sure what was causing this modulation.
    • However, this is inconsistent with a calculation, see Attachment #2. I assumed modulation depth of 90 mrad, round-trip loss of 100 ppm, and Titm = 1.094%, Tetm = 4.579%, numbers I pulled from the core-optics wiki page.
    • Not sure what effect I've missed out on here - to get the model to match the measurement, I have to either assume a higher cavity finesse, or a much higher round-trip loss (5000ppm), both of which seem implausible.

The main motivation was to get the residual frequency noise of the EX laser when locked to the X arm cavity - but I'll need the V/Hz PDH discriminant to convert the in-loop error signal to frequency units. The idea was to look at the PDH error signal on a scope and match up the horn-to-horn voltage with a model to back out said discriminant, but I'll have to double check my model for errors now given the large mismatch I observe in reflected power.

Attachment 1: IMG_7393.JPG  1.210 MB  Uploaded Mon Apr 8 19:46:36 2019  | Show | Hide all | Show all
Attachment 2: greenModeMatch.pdf  19 kB  Uploaded Mon Apr 8 19:46:51 2019  | Hide | Hide all | Show all
greenModeMatch.pdf
ELOG V3.1.3-